1.find the roots of the equation: z^4+4=0 and z^4-4=0.

2.additional exercises for practice are given below.find the roots of:

2.1.z^8-16=0.

2.2.z^8+16=0.



1
Expert's answer
2021-07-06T16:10:08-0400

"z^4+4=0\\
z^4=-4\\
z^4=4(\cos(\pi)+i\sin(\pi))\\
z^4=4(\cos(2\pi{n}+\pi)+i\sin(2\pi{n}+\pi))\\
z=4^\frac{1}{4}(\cos(\frac{2\pi{n}+\pi}{4})+\\i\sin(\frac{2\pi{n}+\pi)}{4})\\
n=0,1,2,3\\
z_o=4^\frac{1}{4}(\cos(\frac{0+\pi}{4})+\\i\sin(\frac{0+\pi)}{4})\\
z_o=4^\frac{1}{4}(\cos(\frac{\pi}{4})+i\sin(\frac{\pi}{4}))\\
z_1=4^\frac{1}{4}(\cos(\frac{3\pi}{4})+i\sin(\frac{3\pi}{4}))\\
z_2=4^\frac{1}{4}(\cos(\frac{5\pi}{4})+i\sin(\frac{5\pi}{4}))\\
z_3=4^\frac{1}{4}(\cos(\frac{7\pi}{4})+i\sin(\frac{7\pi}{4}))\\"

To solve "z^4-4=0"

"z^4=4\\
z^4 =4(cos(0)+isin(0))\\
z^4=4(\cos(2\pi{n})+i\sin(2\pi{n})\\
z=4^\frac{1}{4}(\cos(\frac{2\pi{n}}{4})+\\i\sin(\frac{2\pi{n})}{4})\\
z_o=4^\frac{1}{4}(\cos(0)+i\sin(0))\\
z_o=4^\frac{1}{4}\\
z_1=4^\frac{1}{4}(\cos(\frac{\pi}{2})+i\sin(\frac{\pi}{2}))\\
z_2=4^\frac{1}{4}(\cos(\pi)+i\sin(\pi))\\
z_3=4^\frac{1}{4}(\cos(\frac{3\pi}{2})+i\sin(\frac{3\pi}{2}))\\"

To solve "z^4-16=0"

"z^4=16\\
z^4 =16(cos(0)+isin(0))\\
z^4=16(\cos(2\pi{n})+i\sin(2\pi{n})\\
z=16^\frac{1}{4}(\cos(\frac{2\pi{n}}{4})+\\i\sin(\frac{2\pi{n})}{4})\\
z_o=16^\frac{1}{4}(\cos(0)+i\sin(0))\\
z_o=16^\frac{1}{4}=2\\
z_1=2(\cos(\frac{\pi}{2})+i\sin(\frac{\pi}{2}))\\
z_2=2(\cos(\pi)+i\sin(\pi))\\
z_3=2(\cos(\frac{3\pi}{2})+i\sin(\frac{3\pi}{2}))\\"

To solve "z^4+16=0"

"z^4=-16\\
z^4=16(\cos(2\pi{n}+\pi)+i\sin(2\pi{n}+\pi))\\
z=16^\frac{1}{4}(\cos(\frac{2\pi{n}+\pi}{4})+\\i\sin(\frac{2\pi{n}+\pi)}{4})\\
n=0,1,2,3\\
z_o=16^\frac{1}{4}(\cos(\frac{0+\pi}{4})+\\i\sin(\frac{0+\pi)}{4})\\
z_o=2(\cos(\frac{\pi}{4})+i\sin(\frac{\pi}{4}))=-2\\
z_1=2(\cos(\frac{3\pi}{4})+i\sin(\frac{3\pi}{4}))\\
z_2=2(\cos(\frac{5\pi}{4})+i\sin(\frac{5\pi}{4}))\\
z_3=2(\cos(\frac{7\pi}{4})+i\sin(\frac{7\pi}{4}))\\"